### v1

@@ 1 Introduction
To enable disentanglement learning and regression-based 3D reconstruction of animals, a large-scale dataset is essential. Collecting real-world data is challenging, primarily due to the scarcity of real-world animal data, and the high costs associated with data collection, unlike the relative abundance of available human data [17]. To address this challenge, we develop a synthetic pipeline, DessiePIPE, capable of generating realistic animal images on-the-fly, to support regression-based 3D animal reconstruction while encouraging disentanglement learning. DessiePIPE utilizes a newly published horse motion capture dataset [39] and the hSMAL model [38], rendering horse images, in various poses (standing, walking, trotting, cantering, eating, bending neck, sitting, rearing), shapes, and species-related textures, with randomized orientations and backgrounds. Additionally, we introduce a novel technique, to create image tuples with controllable variations, allowing the study of disentangled features by producing pairs of images that vary one feature while maintaining others (e.g. different appearances with the same pose and global rotation).
@@ 2 Related Work
Given the challenges of learning animal models from 3D captured data, initial research on animal modeling focused on creating 3D models from images. A pioneering effort by Cashman and Fitzgibbon demonstrates the reconstruction of dolphins and bears from images [9]. Later, Zuffi et al. introduce SMAL [84], the animal equivalent of SMPL, learned from 3D scans of a limited set of toys representing different quadrupeds. Based on the SMAL model, Zuffi et al. [83] introduce an end-to-end regression network to reconstruct zebras, marking an important step towards more detailed and accurate animal 3D model-based monocular reconstruction. The 3D capture of dogs has seen innovation, with Kearney et al. [27] utilizing RGBD images for reconstruction, while Biggs et al. [3, 4] achieve pose and shape estimation from solely image data. Rüegg et al. [65, 64] enhance accuracy by incorporate breed-specific information and ground contact data. Bird modeling has also progressed, employing images and a 3D synthetic model [2, 72] to learn shape variation of birds. Li et al. [38] introduce the hSMAL model, a horse-specific SMAL model that facilitates the recovery of 3D poses from images and aids in lameness detection in horses [45], showing the potential of these technologies for biological and medical applications. VAREN [85] builds upon hSMAL, leveraging real horse scan data.
Many existing regression methods [25, 33, 31, 41, 43, 83, 3] encode images as a single feature vector to directly predict camera, model shape and pose parameters. Notably, some methods that take a different approach. PARE [57] employs part-guided attention from one module alongside 3D features from another to better understand occluded body parts. HybrIK [40] leverages separate modules to generate 3D heatmaps and predict shape and twist angles for inverse kinematics. Pavlakos et al. [57] estimate silhouette and heatmap from images to infer model shape and pose, showing that disentangling shape and pose enhances 3D prediction stability and accuracy. Recent works on dogs, Barc [65] and Bite [64], employ separate pose and shape modules. Unlike these approaches, our work proposes to learn three separate feature spaces for shape, pose, and global information (rotation, translation and camera) prediction in a disentangled fashion.
@@
Disentanglement aims to figure out the underlying latent representations that control different data properties. It has be applied to both 2D images [11, 21, 48] and 3D meshes [82]. Lorenz et al. [48] disentangle shape and appearance from 2D images by learning parts consistency over category instances, leveraging the equivariance and invariance constraints on the image representation to extract part-based information. Zhou et al. [82] disentangle shape and pose features from registered meshes by using self-consistency and cross-consistency constraints in an unsupervised setting. Xu et al. [22] extract disentangled 3D attributes only from 2D images to manipulate the 3D output, based on PiFu, which decodes image features from the backbone network into 3D implicit representations of humans. The image features encode three attributes, which are disentangled by the designed frameworks. Zhao et al. [81] disentangle pose and view factors from 2D human poses by maximizing mutual information of the same pose from different views in a contrastive learning manner. While these works focus on 2D pose estimation, 3D mesh generation or 3D mesh manipulation tasks, none address monocular model-based 3D reconstruction.
@@ 3.4 Loss
$L_{SIL}=\omega_{S}smoothL1(\tilde{S},S_{gt}),\ \tilde{S}=\Pi(\mathbf{v},\Omega)~{},$ (4)
@@ 4.3 Experiments
Figure 8: Failure cases. Novel shapes not captured (top). Two subjects not captured (bottom).
@@ 5 Conclusion
5.1 Limitations
Currently, Dessie operates on cropped images with a single subject, which poses challenges if this setting is violated (Fig. 8, bottom). To overcome the limited shape expressiveness of the hSMAL model, which generalizes poorly to shapes very different from those in the training set (Fig. 8, top), the latest horse shape model, VAREN [85] can be used. Finally, while model-based approaches can resolve ambiguities deriving from monocular settings, they require learning a shape and pose prior, which can be challenging for the animal species of interest.
Acknowledgements. We thank Qingwen Zhang, Alfredo Reichlin, and Giovanni Luca Marchetti for the fruitful discussions. Silvia Zuffi is supported by PNRR FAIR Future AI Research (PE00000013), Spoke 8 Pervasive AI (CUP H97G22000210007) under the NRRP MUR program by NextGenerationEU. Elin Hernlund is supported by an SLU career grant no. 57122058, and the Beijer foundation. The computations were enabled by the supercomputing resource Berzelius provided by the National Supercomputer Centre at Linköping University and the Knut and Alice Wallenberg Foundation, Sweden.

### v2

@@ 1 Introduction
To enable disentanglement learning and regression-based 3D reconstruction of animals, a large-scale dataset is essential. Collecting real-world data is challenging, primarily due to the scarcity of real-world animal data, and the high costs associated with data collection, unlike the relative abundance of available human data [17]. To address this challenge, we develop a synthetic pipeline, DessiePIPE, capable of generating realistic animal images on the fly, to support regression-based 3D animal reconstruction while encouraging disentanglement learning. DessiePIPE utilizes a newly published horse motion capture dataset [39] and the hSMAL model [38], rendering horse images, in various poses (standing, walking, trotting, cantering, eating, bending neck, sitting, rearing), shapes, and species-related textures, with randomized orientations and backgrounds. Additionally, we introduce a novel technique, to create image tuples with controllable variations, allowing the study of disentangled features by producing pairs of images that vary one feature while maintaining others (e.g. different appearances with the same pose and global rotation).
@@ 2 Related Work
Given the challenges of learning animal models from 3D captured data, initial research on animal modeling focused on creating 3D models from images. A pioneering effort by Cashman and Fitzgibbon demonstrates the reconstruction of dolphins and bears from images [9]. Later, Zuffi et al. introduce SMAL [84], the animal equivalent of SMPL, learned from 3D scans of a limited set of toys representing different quadrupeds. Based on the SMAL model, Zuffi et al. [83] introduce an end-to-end regression network to reconstruct zebras, marking an important step towards more detailed and accurate animal 3D model-based monocular reconstruction. The 3D capture of dogs has seen innovation, with Kearney et al. [27] utilizing RGBD images for reconstruction, while Biggs et al. [3, 4] achieve pose and shape estimation from solely image data. Rüegg et al. [65, 64] enhance accuracy by incorporating breed-specific information and ground contact data. Bird modeling has also progressed, employing images and a 3D synthetic model [2, 72] to learn shape variation of birds. Li et al. [38] introduce the hSMAL model, a horse-specific SMAL model that facilitates the recovery of 3D poses from images and aids in lameness detection in horses [45], showing the potential of these technologies for biological and medical applications. VAREN [85] builds upon hSMAL, leveraging real horse scan data.
Many existing regression methods [25, 33, 31, 41, 43, 83, 3] encode images as a single feature vector to directly predict camera, model shape and pose parameters. Notably, some methods take a different approach. PARE [30] employs part-guided attention from one module alongside 3D features from another to better understand occluded body parts. HybrIK [40] leverages separate modules to generate 3D heatmaps and predict shape and twist angles for inverse kinematics. Pavlakos et al. [57] estimate silhouette and heatmap from images to infer model shape and pose, showing that disentangling shape and pose enhances 3D prediction stability and accuracy. Recent works on dogs, Barc [65] and Bite [64], employ separate pose and shape modules. Unlike these approaches, our work proposes to learn three separate feature spaces for shape, pose, and global information (rotation, translation and camera) prediction in a disentangled fashion.
@@
Disentanglement aims to figure out the underlying latent representations that control different data properties. It has be applied to both 2D images [11, 21, 48] and 3D meshes [82]. Lorenz et al. [48] disentangle shape and appearance from 2D images by learning parts consistency over category instances, leveraging the equivariance and invariance constraints on the image representation to extract part-based information. Zhou et al. [82] disentangle shape and pose features from registered meshes by using self-consistency and cross-consistency constraints in an unsupervised setting. Hu et al. [22] extract disentangled 3D attributes only from 2D images to manipulate the 3D output, based on PiFu, which decodes image features from the backbone network into 3D implicit representations of humans. The image features encode three attributes, which are disentangled by the designed frameworks. Zhao et al. [81] disentangle pose and view factors from 2D human poses by maximizing mutual information of the same pose from different views in a contrastive learning manner. While these works focus on 2D pose estimation, 3D mesh generation or 3D mesh manipulation tasks, none address monocular model-based 3D reconstruction.
@@ 3.4 Loss
$L_{SIL}=\omega_{S}smoothL1(\tilde{S},S_{gt}),\ \tilde{S}=\Pi(\mathbf{v},\Omega),$ (4)
@@ 5 Conclusion
Limitations.
Currently, Dessie operates on cropped images with a single subject, which poses challenges if this setting is violated (Fig. 8, right). To overcome the limited shape expressiveness of the hSMAL model, which generalizes poorly to shapes very different from those in the training set (Fig. 8, left), the latest horse shape model, VAREN [85] can be used. Finally, while model-based approaches can resolve ambiguities deriving from monocular settings, they require learning a shape and pose prior, which can be challenging for the animal species of interest.
Acknowledgements.
We thank Qingwen Zhang, Alfredo Reichlin, and Giovanni Luca Marchetti for the fruitful discussions. Silvia Zuffi is supported by PNRR FAIR Future AI Research (PE00000013), Spoke 8 Pervasive AI (CUP H97G22000210007) under the NRRP MUR program by NextGenerationEU. Elin Hernlund is supported by an SLU career grant no. 57122058, and the Beijer foundation. The computations were enabled by the supercomputing resource Berzelius provided by the National Supercomputer Centre at Linköping University and the Knut and Alice Wallenberg Foundation, Sweden.
Figure 8: Failure cases. Left: Novel shapes. Right: Two subjects.